OpenStudy (anonymous):

how to graph this? 3x+2y=6

OpenStudy (tkhunny):

Trust me on this. Divide everything by 6.

OpenStudy (anonymous):

so how to graph it tho?

OpenStudy (tkhunny):

I am answering that with my instruction. Did you do it?

OpenStudy (anonymous):

yeah

OpenStudy (tkhunny):

What did you get?

OpenStudy (anonymous):

0

OpenStudy (tkhunny):

?? How can you get that? 3x+2y=6 Divide by 6 \(\dfrac{x}{2} + \dfrac{y}{3} = 1\)

OpenStudy (anonymous):

I did not get It(((

OpenStudy (tkhunny):

I asked you if you did the division by 6. You said "yeah". That means you did not do it? Very odd. Anyway, do you see how the division by 6 works? 3x/6 = x/2 2y/6 = y/3 6/6 = 1

OpenStudy (anonymous):

yeah I see

OpenStudy (tkhunny):

This gives the little-known "Intercept Form". x/2 suggests an x-intercept of (2,0) y/3 suggests an y-intercept of (0,3) Plot those two points, connect the dots, and you are done!

OpenStudy (anonymous):

thanks